How do you factor 8x^2-28x-60
Found 3 solutions by stanbon, dabanfield, richwmiller:
Answer by stanbon(75887)

factor 8x^2-28x-60
---
= 4(2x^2 - 7x - 15)
---
= 4(2x^2 -10x + 3x - 15)
--
= 4(2x(x-5)+3(x-5))
---
= 4(x-5)(2x+3)
==================
Cheers,
Stan H.

Answer by dabanfield(803)

How do you factor 8x^2-28x-60
Divide both sides by 4 just to simplify things:
2x^2 - 7x - 15
(2x - ?) * (x - ?)
Factors of 15 are 3, 5, 15 and 1
(2x + 3) * (x - 5)
Make sense?

Answer by richwmiller(17219)

Quietly at night while eating cherry pie and ice cream or in broad daylight while measuring a sausage and mushroom pizza with a beverage of your choice.
other wise
4 (x-5) (2 x+3)
Let me know your preferred method.
You buy; I eat.